# Cody D. Schimming, C. J. O. Reichhardt, and C. Reichhardt

### v1

@@ 3.2 Two $-1/2$ Defects
In Fig. 6(a) we plot $r_{C}/\ell_{h}$ versus $\alpha/\ell_{h}^{2}$ for $\ell_{h}=10$. Here $r_{C}\to\infty$ as $\alpha\to 0$ as expected, since passive ±$\pm 1/2$ defect pairs will always annihilate. As shown by the log-log plot in Fig. 6(b), we find that $r_{C}\sim\alpha^{-1/2}$, a scaling that has been observed for $\langle r_{d}\rangle$ in continuum simulations 19, 51. In some sense, this scaling is expected because the active length scale $\ell_{a}\propto\alpha^{-1/2}$; however, a naive approximation would lead to the expectation that $dr/dt\sim v-1/r$, where $v$ is the speed of the $+1/2$ defect, and hence that $r_{C}\sim\alpha^{-1}$. Only by including all of the hydrodynamic effects can the $\alpha^{-1/2}$ scaling be predicted. Interestingly, we find that the screening length $\ell_{h}$ plays an important role in this scaling. In Fig. 6(b) we also plot $r_{C}/\ell_{h}$ versus $\alpha/\ell_{h}^{2}$ for $\ell_{h}\in\{1,2,5\}$. For small screening lengths and small activity, $r_{C}\sim\alpha^{-1}$, confirming the naive approximation. In this regime, the self propulsion is the dominant interaction term, and only at higher activities is the $\alpha^{-1/2}$ scaling restored. The crossover between regimes as a function of $\alpha$ occurs for higher relative activities at smaller screening lengths. Thus, we predict that active nematic experiments and simulations performed in the “dry” limit (smaller screening length) for relatively small activities will measure $\langle r_{d}\rangle\sim\alpha^{-1}$.
@@
We find that there is a critical global phase, $\varphi_{C}$, depending on initial separation and activity, above which $+1/2$ defects will deflect and not annihilate. To study this systematically, we compute trajectories of ±$\pm 1/2$ defect pairs with varying initial distances and orientations. We consider the two defects “annihilated” if they come within one nematic correlation length of each other. In our normalized units this occurs when $r<1$. In Fig. 7(b) we plot the critical angle $\varphi_{C}$ versus $\alpha/\ell_{h}^{2}$ for $\ell_{h}=10$ at various initial distances, and in Fig. 7(c) we plot $\varphi_{C}$ versus $r_{0}/\ell_{h}$ for various activities. The insets of these figures show that $\varphi_{C}\sim r_{0}^{-2}\alpha^{-1}$, or $\varphi_{C}\sim(\ell_{a}/r_{0})^{2}$. We can also map this critical angle to a critical impact parameter $b_{C}$ by rotating our axes by $2\varphi_{C}$ (so that $\varphi_{0}\to 0$). Then, $b_{C}=r_{0}\sin 2\varphi_{C}$ so that $b_{C}\approx 2\ell_{a}^{2}/r_{0}$. We sketch the rotation leading to this mapping in Fig. 8.
@@ 4 Multiple Defects in Circular Confinement
where the polar angles of each defect $\theta_{i}$ are subtracted so that $\varphi_{0}$ gives the director angle offset at $R$. Equation (31) describes a director field which will always have a fixed director at $R$. If the total topological charge of the system is 00, then $\phi(r=R)=\varphi_{0}$, a constant. The total topological charge does not have to be zero, however. Equation (31) will give the lowest energy director regardless of the total topological charge. We can then obtain the orientation of each defect using $\varphi_{j}=\lim_{\epsilon\to 0}\phi(x_{j}+\epsilon,y_{j})$.
@@
In confined systems, it is typical to assume no-slip boundary conditions for the fluid velocity. However, to our knowledge, no analytical solution to Eq. (11) is available for $\mathbf{u}(R)=0$ boundary conditions. Thus, here we neglect the effect of the boundaries on the flow solution, and limit our considerations to the case of small screening length, namely $\ell_{h}=1$. In this parameter regime, it was shown in Ref. 39 that the flow solutions for a single defect are only weakly perturbed by no slip boundary conditions when the analytical solution for an infinite domain was compared to bounded numerical solutions. Finally, we assume that defects in the system do not feel the flow solutions from the image charges but experience only the Coulomb force from elastic forces. For the parameters that we explore, this assumption only results in small quantitative changes to the defect motion. In what follows, we integrate the equations of motion for each defect using a forward Euler method with time step $\Delta t=0.001$ and initialize the system with randomly generated positions of the defects.
We first examine multiple $+1/2$ defects in confinement. For two $+1/2$ defects, the results of Sec. 3 apply and the defects will rotate around one another if $\varphi_{0}\neq 0,\pi/2$. If $\varphi_{0}=0,\pi/2$, the defects will find stable positions with $\mathbf{v}=0$. For three $+1/2$ defects, we find that the defects tend to braid one another in a manner similar to that seen in recent experiments 34. In Fig. 9(a,b) we show example trajectories for three defects with system radii $R=3$ and $R=5$, respectively. For the smaller system size, the defects maintain tighter “figure eight” trajectories, as seen in the experiments, until eventually settling into the stable figure eight trajectory shown by the dashed line in Fig. 9(a). For the larger system size the defects do not follow perfect braiding trajectories, but instead traverse figure eight pathways bounded by outer and inner figure eight loops. Here we do not observe the same stabilization of the trajectories that appears in the more confined system. Movies of the integrated analytical model are available in the Supplemental Material (1,2). The ability of our model to predict the braiding behavior seen in experiments highlights the importance of orientational coupling in the active nematic. Such behavior could not be reproduced with a simpler active Brownian particle model where only Coulomb interactions are present.
@@ 5 Discussion
As we have demonstrated, our analytical model qualitatively reproduces features of active nematic defect interactions observed in both continuum simulations and experiments. One of the primary advantages of the analytical model is the ease with which the effect of different parameters may be investigated. For example, in future studies, the model can be used to probe the effect of screening length and tumbling parameter. Another advantage is the modular nature of the theory. One may, for example, investigate the effect of elastic interactions by changing the coefficient for the Coulomb force term; or one could impose different flow profiles by adding the appropriate terms to the velocity. Our analytical model achieves its greatest accuracy under the separation of length scales $a<\langle r_{d}\rangle<\ell_{a}$, where $a$ is the core size, $\langle r_{d}\rangle$ is the average separation of defects, and $\ell_{a}$ is the active length scale. Our linear perturbation approximation, which allows us to treat the flows of individual defects separately, breaks down when the cores of defects overlap. Additionally, we have neglected the effect of topologically trivial structures such as bend walls, which also generate flows in the system 38, on the motion of defects. When $\langle r_{d}\rangle>a$, the effect of defect core overlap will be minimal, and when $\ell_{a}>\langle r_{d}\rangle$, structures such as bend walls will be quickly relaxed due to elastic forces.
@@
We make the approximation that the director field is always given by the minimal energy configuration. It has been shown in experiments that this is not always the case for active nematics 52, 37, and induced torques and transverse velocities may play an important role in the motion of active nematic defects 43, 44. This presents an opportunity for a future extension of the analytical model, since it has been shown that the defect velocity formalism can capture the transverse motion induced by these torques 15. However, to our knowledge, the flows induced by the relative twisting of defects in active nematics are not yet well understood. If approximate equations for this can be derived, then the orientation of topological defects in active nematics may be treated independently of their positions and a true particle model could be formulated. Accurately treating the orientation of defects independently would allow for many additional features to be included in the model such as defect nucleation, or orientational noise. It would also be possible to include periodic boundary conditions, which cannot be treated in the present model due to instantaneous jumps in the orientation when defects cross the periodic boundary. In each case, it would be possible to maintain the orientational coupling that we have shown is important when describing active nematic defect motion.
@@
An interesting extension of the methodology presented here would be to treat line defects in three-dimensional active nematics. Although it has been shown that the defect velocity formalism we employ here can be generalized to line defects in three dimensions 15, there has been much less work in understanding defect dynamics in three dimensional active nematics both computationally 53, 54, and, especially, experimentally 55. Recent theoretical work has described some aspects of defect loop motion 56, 57, and it will be interesting to compare and combine the results of the methodologies used here to obtain a clearer picture of defect motion in active nematics.
@@ 6 Summary
We have presented an analytical model for active nematic defect motion based on recently derived theoretical techniques and solutions to the Stokes equation for active nematic defects. The model incorporates both the flow-induced long range interactions between defects and also the crucial orientational coupling between the defects. We demonstrate that our model provides analytical predictions for the stable interaction radius adopted by bound states of pairs of $+1/2$ defects, the effective attraction that appears between pairs of $-1/2$ defects, and the way in which the critical unbinding length between a $+1/2$ and $-1/2$ defect pair scales with activity. The model can also capture the dependence of the scattering trajectory of a $+1/2$ defect from a $-1/2$ defect on activity. It can be applied to the braiding motion of multiple confined $+1/2$ defects as well as the coarsening behavior that occurs for a mixture of equal numbers of $+1/2$ and $-1/2$ defects under confinement. Our work provides an alternative picture to studying defect dynamics in active nematics and lays the groundwork for the development of a full-fledged particle model. It complements the traditional analysis of active nematics through continuum models and will serve as a useful tool for future studies.
@@
Here we present more detailed calculations leading to the expressions in Eqs. (5), (6), and (10). For completeness, we start by briefly reviewing the arguments leading to Eq. (2), though more detailed derivations may be found in Refs. 11, 15.
@@
To get analytical approximations for the velocity, we must approximate $\mathbf{Q}$ near defect cores. We assume a linear core structure 12:
$\mathbf{Q}(x,y)\approx\frac{x}{2}\left[\mathbf{\tilde{n}}_{0}\otimes\mathbf{\tilde{n}}_{0}-\mathbf{\tilde{n}}_{1}\otimes\mathbf{\tilde{n}}_{1}\right]\\ +\frac{y}{2}\left[\mathbf{\tilde{n}}_{0}\otimes\mathbf{\tilde{n}}_{1}+\mathbf{\tilde{n}}_{1}\otimes\mathbf{\tilde{n}}_{0}\right]$ (36)
@@
The Coulomb velocity in Eq. (5) may be derived by substituting $\partial_{t}Q_{\mu\alpha}=-\delta F/\delta Q_{\mu\alpha}=(1/k)\nabla^{2}Q_{\mu\alpha}$ into Eq. (2). Using Eq. (36) to compute the spatial derivative and setting $x=y=0$ gives Eq. (5). Given a fluid flow $\mathbf{u}(\mathbf{r}=0)$, substituting $\partial_{t}Q_{\mu\alpha}=-(\mathbf{u}\cdot\nabla)\mathbf{Q}_{\mu\alpha}$ into Eq. (2) gives the advection component of the velocity, Eq. (6), after using Eq. (36) to compute the gradient of $\mathbf{Q}$. To get the shear component of the defect velocity, Eq. (10), we first note that upon substituting $\partial_{t}Q_{\mu\alpha}=S_{\mu\alpha}$ into Eq. (2), most of the terms are zero when $x=y=0$ since $\mathbf{Q}=0$ at the core of the defect. The only term that contributes to the velocity is $\partial_{t}Q_{\mu\alpha}=\lambda E_{\mu\alpha}$, and substitution of this into Eq. (2) yields Eq. (10).
@@
Here we give the functional form of the distance dependent part of the flow and shear flow vectors for a single ±$\pm 1/2$ defect. The flow solutions are equivalent to those derived in Ref. 39, though we have rewritten them to decompose them into radial and angular parts. The radial parts written here are infinite series which are tabulated to find the results in the main text.
@@
For all simulations we set $A=-0.5$, so at equilibrium the liquid crystal is in the nematic phase with $S_{N}=\sqrt{2}$. To numerically solve the dynamical equations, Eq. (3) is discretized in time with an implicit backward Euler scheme and a Newton-Raphson sub-iteration to solve the resulting nonlinear equations. Both Eqs. (3) and (11) are discretized in space on either a square domain or a circular domain and we solve the weak form of partial differential equations using the MATLAB/C++ package FELICITY 58. At each time step, given a configuration of $\mathbf{Q}$ we first solve Eq. (11) for $\mathbf{u}$. We then use this solution to compute the next time step for $\mathbf{Q}$. For all simulations we use a time-step of $\Delta t=0.1$.
@@
For the continuum simulation where we compare defect coarsening, shown in Fig. 10, we use Dirichlet boundary conditions with $\phi(R)=0$, requiring a total topological charge of zero. To find the number of defects as a function of time, we integrate the magnitude of $D$, given in Eq. (34) 59, 13:

### v2

@@ 3.2 Two $-1/2$ Defects
In Fig. 6(a) we plot $r_{C}/\ell_{h}$ versus $\alpha/\ell_{h}^{2}$ for $\ell_{h}=10$. Here $r_{C}\to\infty$ as $\alpha\to 0$ as expected, since passive ±$\pm 1/2$ defect pairs will always annihilate. As shown by the log-log plot in Fig. 6(b), we find that $r_{C}\sim\alpha^{-1/2}$, a scaling that has been observed for $\langle r_{d}\rangle$ in continuum simulations 20, 64. In some sense, this scaling is expected because the active length scale (the length scale that arises due to the competition between active and elastic forces 18) $\ell_{a}\propto\alpha^{-1/2}$; however, a naive approximation would lead to the expectation that $dr/dt\sim v-1/r$, where $v$ is the speed of the $+1/2$ defect, and hence that $r_{C}\sim\alpha^{-1}$. Only by including all of the hydrodynamic effects can the $\alpha^{-1/2}$ scaling be predicted. Interestingly, we find that the screening length $\ell_{h}$ plays an important role in this scaling. In Fig. 6(b) we also plot $r_{C}/\ell_{h}$ versus $\alpha/\ell_{h}^{2}$ for $\ell_{h}\in\{1,2,5\}$. For small screening lengths and small activity, $r_{C}\sim\alpha^{-1}$, confirming the naive approximation. In this regime, the self propulsion is the dominant interaction term, and only at higher activities is the $\alpha^{-1/2}$ scaling restored. The crossover between regimes as a function of $\alpha$ occurs for higher relative activities at smaller screening lengths. Thus, we predict that active nematic experiments and simulations performed in the “dry” limit (smaller screening length) for relatively small activities will measure $\langle r_{d}\rangle\sim\alpha^{-1}$.
To gain some analytical insight we use the approximation of $\beta_{1}$ in the limit of $r_{C}<<\ell_{h}$ [see Fig. 2(a) and the Appendix]. Setting $\beta_{1}(r_{C}/2\ell_{h})=(\pi^{2}/4)/(\pi/4+r_{C}/2\ell_{h})$ in Eq. (30) and neglecting the terms which are small in this limit gives
$r_{C}\approx\frac{1+\sqrt{1+\ell_{h}^{2}(\pi^{2}/4)(\alpha/\ell_{h}^{2})}}{\ell_{h}(\pi/4)(\alpha/\ell_{h}^{2})}$ (31)
where we have written the right hand side explicitly in terms of $\alpha/\ell_{h}^{2}$ to match the scaling in Fig. 6. From this approximation, we see that if either $\ell_{h}$ or $\alpha$ is large then we recover $r_{C}\sim(\alpha/\ell_{h}^{2})^{-1/2}$, while if both $\ell_{h}$ and $\alpha$ are small $r_{C}\sim(\alpha/\ell_{h}^{2})^{-1}$. Thus, both limits are able to be captured by this approximation, and we emphasize that the full model can interpolate between these two regimes.
@@
We find that, if defects are initially oriented along the $x$-axis, there is a critical global phase, $\varphi_{C}$, depending on initial separation and activity, above which $+1/2$ defects will deflect and not annihilate. To study this systematically, we compute trajectories of ±$\pm 1/2$ defect pairs with varying initial distances and orientations. We consider the two defects “annihilated” if they come within one nematic correlation length of each other. In our normalized units this occurs when $r<1$. In Fig. 7(b) we plot the critical angle $\varphi_{C}$ versus $\alpha/\ell_{h}^{2}$ for $\ell_{h}=10$ at various initial distances, and in Fig. 7(c) we plot $\varphi_{C}$ versus $r_{0}/\ell_{h}$ for various activities. The insets of these figures show that $\varphi_{C}\sim r_{0}^{-2}\alpha^{-1}$, or $\varphi_{C}\sim(\ell_{a}/r_{0})^{2}$. We can also map this critical angle to a critical impact parameter $b_{C}$ by rotating our axes by $2\varphi_{C}$ (so that $\varphi_{0}\to 0$). Then, $b_{C}=r_{0}\sin 2\varphi_{C}$ so that $b_{C}\approx 2\ell_{a}^{2}/r_{0}$. We sketch the rotation leading to this mapping in Fig. 8.
Our analysis of the model’s predictions for the critical unbinding lengths and angles of ±$\pm 1/2$ annihilation do not assume a particular method of nucleation. One future extension of the model would be to implement specific nucleation methods. However, once the defects have nucleated, our analysis should apply to situations where the ±$\pm 1/2$ pair are well separated or screened from other defects in the system. Therefore, we expect our analysis to apply to defects nucleated from the usual bend instability 18, as well as defects nucleated via recently observed disorder-order transitions in phototactic active nematics 65.
@@ 4 Multiple Defects in Circular Confinement
where the polar angles of each defect $\theta_{i}$ are subtracted so that $\varphi_{0}$ gives the director angle offset at $R$. Equation (32) describes a director field which will always have a fixed director at $R$. If the total topological charge of the system is zero, then $\phi(r=R)=\varphi_{0}$, a constant. The total topological charge does not have to be zero, however. Equation (32) will give the lowest energy director regardless of the total topological charge. We can then obtain the orientation of each defect using $\varphi_{j}=\lim_{\epsilon\to 0}\phi(x_{j}+\epsilon,y_{j})$.
@@
In confined systems, it is typical to assume no-slip boundary conditions for the fluid velocity. However, to our knowledge, no analytical solution to Eq. (11) is available for $\mathbf{u}(R)=0$ boundary conditions. Thus, here we neglect the effect of the boundaries on the flow solution, and limit our considerations to the case of small screening length, namely $\ell_{h}=1$. In this parameter regime, it was shown in Ref. 43 that the flow solutions for a single defect are only weakly perturbed by no slip boundary conditions when the analytical solution for an infinite domain was compared to bounded numerical solutions. Finally, we assume that defects in the system do not feel the flow solutions from the image charges but experience only the Coulomb force from elastic forces. For the parameters that we explore, this assumption only results in small quantitative changes to the defect motion. In what follows, we integrate the equations of motion for each defect using a forward Euler method with time step $\Delta t=0.001$ and initialize the system with randomly generated positions of the defects.
We first examine multiple $+1/2$ defects in confinement. For two $+1/2$ defects, the results of Sec. 3 apply and the defects will rotate around one another if $\varphi_{0}\neq 0,\pi/2$. If $\varphi_{0}=0,\pi/2$, the defects will find stable positions with $\mathbf{v}=0$. For three $+1/2$ defects, we find that the defects tend to braid one another in a manner similar to that seen in recent experiments 36. In Fig. 9(a,b) we show example trajectories for three defects with system radii $R=3$ and $R=5$, respectively. For the smaller system size, the defects maintain tighter “figure eight” trajectories, as seen in the experiments, until eventually settling into the stable figure eight trajectory shown by the dashed line in Fig. 9(a). For the larger system size the defects do not follow perfect braiding trajectories, but instead traverse figure eight pathways bounded by outer and inner figure eight loops. Here we do not observe the same stabilization of the trajectories that appears in the more confined system. Movies of the integrated analytical model are available in the Supplemental Material (1,2). The ability of our model to predict the braiding behavior seen in experiments highlights the importance of orientational coupling in the active nematic. Such behavior could not be reproduced with a simpler active Brownian particle model where only Coulomb interactions are present.
@@ 5 Discussion
As we have demonstrated, our analytical model qualitatively reproduces features of active nematic defect interactions observed in both continuum simulations and experiments. It is also able to reproduce predictions from previous analytical models of defect motion, such as the mutual rotation of two $+1/2$ defects, and the deflection of $+1/2$ defects by $-1/2$ defects 12, 49. By including a well-defined screening length, the model can interpolate between wet and dry regimes of active nematics. Additionally, because we employ a linear perturbation approximation, the model may be used to investigate the motion of multiple defects interacting with one another.
One of the primary advantages of the analytical model is the ease with which the effect of different parameters may be investigated. For example, in future studies, the model can be used to probe the effect of screening length and tumbling parameter. Another advantage is the modular nature of the theory. One may, for example, investigate the effect of elastic interactions by changing the coefficient for the Coulomb force term; or one could impose different flow profiles by adding the appropriate terms to the velocity. Since the velocity of defects is derived explicitly from topological charge conservation via the methods of Halperin and Mazenko51, 52, 53, 54, we do not need to interpret the results in terms of auxiliary fields38, 39. Having an explicit form for the defect velocities is particularly amenable for computational implementation.
@@
Our analytical model achieves its greatest accuracy under the separation of length scales $a<\langle r_{d}\rangle<\ell_{a}$, where $a$ is the core size, $\langle r_{d}\rangle$ is the average separation of defects, and $\ell_{a}$ is the active length scale. Our linear perturbation approximation, which allows us to treat the flows of individual defects separately, breaks down when the cores of defects overlap. Additionally, we have neglected the effect of topologically trivial structures such as bend walls, which also generate flows in the system 48, on the motion of defects. When $\langle r_{d}\rangle>a$, the effect of defect core overlap will be minimal, and when $\ell_{a}>\langle r_{d}\rangle$, structures such as bend walls will be quickly relaxed due to elastic forces.
@@
We make the approximation that the director field is always given by the minimal energy configuration. It has been shown in experiments that this is not always the case for active nematics 66, 47, and induced torques, transverse velocities, and orientational noise may play an important role in the motion of active nematic defects 58, 9, 24, 11. This presents an opportunity for a future extension of the analytical model, since it has been shown that the defect velocity formalism used here can capture the transverse motion induced by these torques 16. However, to our knowledge, the flows induced by the relative twisting of defects in active nematics are not yet well understood. If approximate equations for this can be derived, then the orientation of topological defects in active nematics may be treated independently of their positions, while still taking into account the intrinsic coupling of the director field and defect orientation, and a true particle model could be formulated using the present theoretical framework. Accurately treating the orientation of defects independently would allow for many additional features to be included in the model such as defect nucleation or orientational noise. It would also be possible to include periodic boundary conditions, which cannot be treated in the present model due to instantaneous jumps in the orientation when defects cross the periodic boundary. In each case, it would be possible to maintain the orientational coupling that we have shown is important when describing active nematic defect motion.
@@
An interesting extension of the methodology presented here would be to treat line defects in three-dimensional active nematics. Although it has been shown that the defect velocity formalism we employ here can be generalized to line defects in three dimensions 16, 67, there has been much less work in understanding defect dynamics in three dimensional active nematics both computationally 68, 69, and, especially, experimentally 70. Recent theoretical work has described some aspects of defect loop motion 71, 72, and it will be interesting to compare and combine the results of the methodologies used here to obtain a clearer picture of defect motion in active nematics.
@@ 6 Summary
We have presented an analytical model for active nematic defect motion based on recently derived theoretical techniques and solutions to the Stokes equation for active nematic defects. Our work builds upon previous efforts to understand defect motion by applying a linear perturbation approximation to the Halperin-Mazenko formalism of defect dynamics. Using this, explicit velocities for an arbitrary number of defects may be predicted. The model incorporates both the flow-induced long range interactions between defects and also the crucial orientational coupling between the defects. We demonstrate that our model provides analytical predictions for the stable interaction radius adopted by bound states of pairs of $+1/2$ defects, the effective attraction that appears between pairs of $-1/2$ defects, and the way in which the critical unbinding length between a $+1/2$ and $-1/2$ defect pair scales with activity. The model can also capture the dependence of the scattering trajectory of a $+1/2$ defect from a $-1/2$ defect on activity. It can be applied to the braiding motion of multiple confined $+1/2$ defects as well as the coarsening behavior that occurs for a mixture of equal numbers of $+1/2$ and $-1/2$ defects under confinement. It will serve as a useful tool for future studies.
@@
Here we present more detailed calculations leading to the expressions in Eqs. (5), (6), and (10). For completeness, we start by briefly reviewing the arguments leading to Eq. (2), though more detailed derivations may be found in Refs. 12, 16 for nematic defects and Refs. 52, 53, 54 for more general systems of defects.
@@
To get analytical approximations for the velocity, we must approximate $\mathbf{Q}$ near defect cores. We assume a linear core structure 13:
$\mathbf{Q}(x,y)\approx\frac{x}{2}\left[\mathbf{\tilde{n}}_{0}\otimes\mathbf{\tilde{n}}_{0}-\mathbf{\tilde{n}}_{1}\otimes\mathbf{\tilde{n}}_{1}\right]+\frac{y}{2}\left[\mathbf{\tilde{n}}_{0}\otimes\mathbf{\tilde{n}}_{1}+\mathbf{\tilde{n}}_{1}\otimes\mathbf{\tilde{n}}_{0}\right]$ (37)
@@
The Coulomb velocity in Eq. (5) may be derived by substituting $\partial_{t}Q_{\mu\alpha}=-\delta F/\delta Q_{\mu\alpha}=(1/k)\nabla^{2}Q_{\mu\alpha}$ into Eq. (2). Using Eq. (37) to compute the spatial derivative and setting $x=y=0$ gives Eq. (5). Given a fluid flow $\mathbf{u}(\mathbf{r}=0)$, substituting $\partial_{t}Q_{\mu\alpha}=-(\mathbf{u}\cdot\nabla)Q_{\mu\alpha}$ into Eq. (2) gives the advection component of the velocity, Eq. (6), after using Eq. (37) to compute the gradient of $\mathbf{Q}$. To get the shear component of the defect velocity, Eq. (10), we first note that upon substituting $\partial_{t}Q_{\mu\alpha}=S_{\mu\alpha}$ into Eq. (2), most of the terms are zero when $x=y=0$ since $\mathbf{Q}=0$ at the core of the defect. The only term that contributes to the velocity is $\partial_{t}Q_{\mu\alpha}=\lambda E_{\mu\alpha}$, and substitution of this into Eq. (2) yields Eq. (10).
@@
Here we give the functional form of the distance dependent part of the flow and shear flow vectors for a single ±$\pm 1/2$ defect. The flow solutions are equivalent to those derived in Ref. 43, though we have rewritten them to decompose them into radial and angular parts. The radial parts written here are infinite series which are tabulated to find the results in the main text.
@@
To obtain the small $x$ approximation of $\beta_{1}$ we compute the first two terms giving $\beta_{1}(x)=\pi-4x+\ldots$ Comparing to a function of the form $f(x)=B/(A+x)$, we may write (for $x<1$) $f(x)=B/A-(B/A^{2})x+\ldots$ Thus, assuming $\beta_{1}(x)$ scales as $f(x)$ for small $x$, we identify $B=\pi^{2}/4$ and $A=\pi/4$, or $\beta_{1}(x)\approx(\pi^{2}/4)/(\pi/4+x)$. The inset of Fig. 2(a) shows empirically that this is a reasonable approximation for small $x$.
@@
For all simulations we set $A=-0.5$, so at equilibrium the liquid crystal is in the nematic phase with $S_{N}=\sqrt{2}$. To numerically solve the dynamical equations, Eq. (3) is discretized in time with an implicit backward Euler scheme and a Newton-Raphson sub-iteration to solve the resulting nonlinear equations. Both Eqs. (3) and (11) are discretized in space on either a square domain or a circular domain and we solve the weak form of partial differential equations using the MATLAB/C++ package FELICITY 73. At each time step, given a configuration of $\mathbf{Q}$ we first solve Eq. (11) for $\mathbf{u}$. We then use this solution to compute the next time step for $\mathbf{Q}$. For all simulations we use a time-step of $\Delta t=0.1$.
@@
For the continuum simulation where we compare defect coarsening, shown in Fig. 10, we use Dirichlet boundary conditions with $\phi(R)=0$, requiring a total topological charge of zero. To find the number of defects as a function of time, we integrate the magnitude of $D$, given in Eq. (35) 74, 14: